OpenStudy (bekkah323):

using y=5x+70 and y=5x+80 form a system of two linear equations in two variables.

OpenStudy (bekkah323):

@skullpatrol

OpenStudy (anonymous):

solve one of the equations for x, then plug your answer into the x value on the other equation and solve for Y

OpenStudy (anonymous):

try bring the 80 to the other side of the = and the making x by itself. then use what x= in that equation and put it into the other one

OpenStudy (bekkah323):

so if i bring the 80 over it would be y-80=5x

OpenStudy (anonymous):

yupp, and since both equations have a "5x" in them, you can replace the "5x" in the other formula and solve for y

OpenStudy (bekkah323):

ok so it would be y=(y-80)+70

OpenStudy (bekkah323):

but then wouldn't my y's cancel out

OpenStudy (anonymous):

you have to solve the other equation for 5x as well before you substitute

OpenStudy (anonymous):

\[5x=y-80\] and \[5x=y-70\]

OpenStudy (bekkah323):

so y-80=5x and y-70=5x

OpenStudy (bekkah323):

then what

OpenStudy (anonymous):

the replace one of the 5x with the equation for the other

OpenStudy (anonymous):

no sorry wrong, been a while since i did this. you have to add the two equations first

OpenStudy (bekkah323):

?

OpenStudy (anonymous):

ok so

OpenStudy (anonymous):

\[0=5x+10-y\]

OpenStudy (anonymous):

70 not 10

OpenStudy (bekkah323):

ok

OpenStudy (bekkah323):

the the other one would be 0=5x+80-y?

OpenStudy (anonymous):

and then plug in for y because its easier. so you get \[0=5x+70-(5x+80)\]

OpenStudy (anonymous):

and you solve for x

OpenStudy (bekkah323):

but then the x cancels out?

OpenStudy (anonymous):

disregard those lines are linear

OpenStudy (bekkah323):

So what does that mean?

OpenStudy (bekkah323):

im supposed to sole the systems using any method we want.

OpenStudy (anonymous):

they are parallel, the question has no answer in the form ax+by=p, cx+dy=q the way a normal linear system would

OpenStudy (bekkah323):

isn't it y=mx+b

OpenStudy (anonymous):

no

OpenStudy (bekkah323):

that the form of the lines i have?

OpenStudy (anonymous):

they want you to prove that these equations are parallel

OpenStudy (anonymous):

so you have to set them equal, \[5x+70=5x+80\] then the 5x cancel out and you are left with 70=80 which we know is not true therefore the two lines never cross and are parallel

OpenStudy (anonymous):

does that make sense to you?

OpenStudy (bekkah323):

yup

OpenStudy (anonymous):

glad I was able to help :)